# Right-handed neutrino dark matter consistent with the generation of baryon number asymmetry

### v1

@@ Abstract
A right-handed neutrino is a promising candidate for dark matter (DM) which has no interaction with nuclei. Since two right-handed neutrinos explain neutrino oscillation data and baryon number asymmetry through seesaw mechanism and leptogenesis based on the out-of-equilibrium decay of the lighter one, the lightest third right-handed neutrino could be free from them and stable to be DM. It suggests a possibility that the right-handed neutrinos could be common key particles for the explanation of three problems in the standard model (SM). We study this possibility by analyzing the structure of a neutrino mass matrix in the scotogenic model for neutrino mass.
@@ 1 Introduction
Neutrino oscillation experiments prove the existence of nonzero neutrino mass [1]. Observations of the abundances of light elements and the cosmic microwave background suggest the existence of the baryon number asymmetry in the Universe [2, 3]. However, both the neutrino mass and the baryon number asymmetry cannot be explained in the standard model (SM). It suggests that the SM is incomplete and has to be extended. If the SM is extended by two right-handed neutrinos which are assumed to have Majorana mass, these two problems are considered to be solved through the so called seesaw mechanism [4] and leptogenesis caused by the out-of-equilibrium decay of the right-handed neutrino [5] as long as their Majorana mass is larger than $O(10^{9})$ GeV [6].
On the other hand, the SM has no candidate of dark matter (DM) whose existence has been proved through various astrophysical and cosmological observations. Direct detection experiments of it severely constrain its mass and scattering cross section with a nucleus, and they exclude a lot of DM candidates from a consistency with the required DM relic abundance [7]. It may suggest a possibility that DM has no interaction with nuclei. In this sense, the right-handed neutrino, which is massive and electrically neutral, could be a good candidate of DM since it has no interaction with nuclei. If we combine this idea with the above solution for the neutrino mass and the baryon number asymmetry, the third right-handed neutrino seems to be a good DM candidate if its stability is guaranteed by some symmetry. This stability requirement may impose to introduce an additional doublet scalar $\eta$ since we have to guarantee the existence of their invariant Yukawa couplings between doublet leptons and right-handed neutrinos under the assumed symmetry. As such an example, we have the scotogenic model [8], which can make the neutrino mass generation possible radiatively around TeV scales. The model seems to be very interesting as a simple extension of the SM since three big problems in the SM might be explained by the right-handed neutrinos simultaneously.
@@
In previous studies aiming to solve these problems including the case where DM is the lightest neutral component of $\eta$, the model is extended by some way, for example, new interaction is introduced to the right-handed neutrinos [11, 12, 13, 14]. In this paper, we propose a solution to these problems within the original scotogenic model without considering any extension of the model.aaaIt is discussed in [15] that $N_{1}$ can be DM consistently with the LFV constraint by considering coannihilation among the right-handed neutrinos without any extension.
The remaining parts are organized as follows. In section 2, we review the model and discuss a neutrino mass matrix derived in it to fix flavor mixing and mass eigenvalues. After that, we study a consistent solution with both the relic DM abundance and the constraint from the lepton flavor violating processes. In section 3, we discuss the consistency of the scenario with successful leptogenesis. We summarize the paper in section 4.
@@ 2 Neutrino model with right-handed neutrino DM
for the vacuum stability. If $\eta$ is assumed to have no vacuum expectation value (VEV), neutrino mass is not produced at tree level and the $Z_{2}$ symmetry is kept as an exact asymmetry of the model.
@@
where $\eta^{\pm}$ stands for charged components, and $\eta_{R}$ and $\eta_{I}$ are real and imaginary components, respectively. In this setting, a lightest right-handed neutrino $N_{1}$ or $\eta_{R}$ can be stable due to the $Z_{2}$ symmetry. We suppose that $N_{1}$ is DM and the baryon number asymmetry is generated by leptogenesis through the out-of-equilibrium decay of a heavier right-handed neutrino. In the following study, mass ordering of these extra particles is assumed to satisfy [16]
@@ 2.1 Neutrino mass and flavor mixing
Although neutrino mass is forbidden at tree level in Lagrangian (1), small neutrino mass is induced through a one-loop diagram caused by the Higgs VEV as
@@
Its diagonalization matrix can be described by using the Yukawa couplings $h_{\alpha k}$ explicitly. When a condition $\sum_{\alpha}(h_{\alpha j}h_{\alpha k})\propto\delta_{jk}$ is satisfied, which corresponds to the condition that matrices $\mu^{j}$ and $\mu^{k}$ are commutable [11], it is diagonalized by a matrix $U$ as $U^{T}{\cal M}_{\nu}U={\rm diag}(m_{1},m_{2},m_{3})$. This $U$ can be expressed as
@@
where $h_{k}$ is a real constant. The absolute value of mass eigenvalues $m_{k}$ is fixed as
@@
where $p_{k}$ is defined as $p_{k}=\left|\sum_{\alpha}U_{\alpha k}^{2}\right|$. Here, it should be noted that $\sum_{\alpha}U_{\alpha k}^{2}=1$ is not satisfied if Yukawa couplings $h_{\alpha k}$ are complex. By using this $U$, the PMNS matrix [17] can be written as $V_{\rm PMNS}=U_{e}^{\dagger}UP$ , where $U_{e}$ is a matrix which is used to diagonalize the charged lepton mass matrix and $P$ is a Majorana phase matrix ${\rm diag}(1,e^{i(\gamma_{2}-\gamma_{1})},e^{i(\gamma_{3}-\gamma_{1})})$. Although $V_{\rm PMNS}$ is unitary in a case where $h_{\alpha k}$ is real, it is not unitary in a case where $h_{\alpha k}$ is complex since $U$ is not a unitary matrix.
@@
$U_{\nu}=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}&c_{12}c_{23}-s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}&-c_{12}s_{23}-s_{12}s_{13}c_{23}&c_{13}c_{23}\\ \end{array}\right),$ (8)
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}e^{i\alpha_{ij}}$. We can represent the neutrino Yukawa couplings $h_{\alpha k}$ and mass eigenvalues $m_{k}$ by using this $U$ as eq. (6) and eq. (7), respectively.
If $h_{\alpha 1}$ has the same structure as $h_{\alpha 2}$ for example, the mass matrix can be diagonalized by $U$ given in eq. (8) but we have a zero mass eigenvalue $m_{1}=0$. On the $m_{2}$ formula in eq. (7), $p_{2}h_{2}^{2}\Lambda_{2}$ is replaced by $p_{1}h_{1}^{2}\Lambda_{1}+p_{2}h_{2}^{2}\Lambda_{2}$. As such a concrete example, neutrino Yukawa couplings may be taken as [14]
@@
The mixing matrix $U$ can be taken as the tribimaximal mixing matrix [18] by fixing the mixing angles as
@@
Since the fault of the tribimaximal mixing matrix as the PMNS matrix could be overcome by considering a nontrivial flavor mixing matrix $U_{e}$ in the charged lepton sector [14], it can become a phenomenologically interesting possibility.
@@
If we use the best fit values from neutrino oscillation data given in [19] for $\sin\theta_{ij}$ in the matrix $U$, we can calculate two effective masses relevant to two observables, that is, the effective masses for the neutrinoless double $\beta$ decay and the $\beta$ decay. They are predicted to be
$m_{\beta\beta}\simeq\left|\sum_{j=1}^{3}U_{ej}^{2}m_{j}\right|=0.0037~{}{\rm eV},\quad m_{\beta}\simeq\sqrt{\sum_{j=1}^{3}|U_{ej}|^{2}m_{j}^{2}}=0.0088~{}{\rm eV}.$ (13)
These results are much smaller than their present bounds [20, 21].
@@
$M_{1}=1.5~{}{\rm TeV},\quad M_{\eta_{R}}\simeq M_{2}\simeq M_{3}\simeq 3~{}{\rm TeV},\quad\lambda_{5}=-10^{-3},$
$h_{1}=6\times 10^{-9},\quad h_{2}=1.11\times 10^{-3},\quad h_{3}=4.15\times 10^{-3},$ (14)
where we assume $\alpha_{12}=\alpha_{13}=0,~{}\alpha_{23}=\frac{\pi}{2}$ and use the best fit values from neutrino oscillation data for $\sin\theta_{ij}$. A smaller $|\lambda_{5}|$ makes $h_{k}$ larger for fixed values of $M_{k}$. Since DM is supposed to be $N_{1}$ here, there is no constraint on a lower bound of the value of $|\lambda_{5}|$ differently from the case where $\eta_{R}$ is DM [10]. We find that the present bounds of LFV like $\mu\rightarrow e\gamma$ [19] can be satisfied for these parameters. We come back to this point later. In the DM study, it may be useful to note that the above values of $h_{2,3}$ are not affected even if $h_{1}$ takes the same order value as $h_{2,3}$ as long as $M_{1}$ is much smaller than $M_{\eta_{R}}$. It is caused by the fact that $\Lambda_{1}$ has a suppression of $O(M_{1}/M_{\eta_{R}})$ compared with $\Lambda_{2,3}$. It allows $N_{1}$ with keV scale mass to have an acceptable relic abundance as DM for the same $h_{2,3}$ given in eq. (14).
Finally, we note an important consequence caused by the feature of this mass matrix addressed above. If we consider leptogenesis based on the out-of-equilibrium decay of the right-handed neutrino $N_{k}$, the model predicts an interesting result. $CP$ asymmetry in this process is known to be caused by the interference between tree and one-loop diagrams, and it should be nonzero at least to make leptogenesis work. Since it is proportional to ${\rm Im}(\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j})^{2}$ for $k\not=j$ as shown in eqs. (21) and (25) later, this condition cannot be satisfied because of $\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j}\propto\delta_{kj}$ for real neutrino Yukawa couplings $h_{\alpha k}$ even if Majorana phases in the right-handed neutrino mass are taken into account. Thus, Yukawa couplings $h_{\alpha k}$ should be complex, and the PMNS matrix cannot be unitary in that case. However, we should note that there is an exceptional case where the $CP$ asymmetry can take a nonzero value even if Yukawa couplings $h_{\alpha k}$ are real. The existence of a zero mass eigenvalue presents such a case as found in an example given in eq. (9) when Majorana phases exist in the right-handed neutrino mass [14]. The PMNS matrix becomes unitary in this case.
@@ 2.2 Relic abundance of right-handed neutrino DM
Right-handed neutrino $N_{1}$ is stable to be DM under the condition (4). However, its abundance is difficult to reach the required value in the freezeout scenario as mentioned in the introduction. It can reach the equilibrium at the temperature higher than its mass through both the $\eta$ decay and the scatterings caused by the neutrino Yukawa couplings $h_{\alpha 1}$ if they have sufficient magnitude. In that case, $h_{\alpha 1}$ has to take a large value of $O(1)$ to reduce its abundance sufficiently through the annihilation. Unfortunately, such a large Yukawa coupling causes LFV processes like $\mu\rightarrow e\gamma$ at an unacceptable level through the one-loop diagram [9]. Since this difficulty is caused by a feature that $N_{1}$ has interaction due to the neutrino Yukawa couplings only, new interaction like $U(1)^{\prime}$ has been introduced to reduce its abundance in many articles [11, 12].
In the present model, since the neutrino Yukawa coupling $h_{1}$ is supposed to be very small as shown in eq. (14), its thermal abundance is originally difficult to reach a value larger than the required one so that the freezeout scenario cannot be applied to the model. However, even in that case, we can consider an alternative possibility such that it is realized through the decay of $\eta_{R}$ following the freezein scenario [22]. $N_{1}$ is produced through the $\eta_{R}$ decay caused by the neutrino Yukawa coupling $h_{1}$. Since $h_{1}$ is supposed to be sufficiently small here, the $N_{1}$ generation through the $\eta_{R}$ decay is expected to be largely suppressed. On the other hand, the next lightest DM candidate $\eta_{R}$ is in the thermal equilibrium through the SM interactions. Decrease of the number density of $\eta_{R}$ is controlled by both the coannihilation [23] to the SM contents and the decay to $N_{1}$. The $N_{1}$ relic abundance is determined by both the $N_{1}$ freezein and the $\eta_{R}$ freezeout depending on their relative effectiveness.
@@
where $H(M_{2})$ represents the Hubble parameter at $T=M_{2}$. Reaction densities $\gamma^{D}(a)$ and $\gamma^{S}(ab)$ are respectively related to thermally averaged decay width $\langle\Gamma_{a}\rangle$ for $a\rightarrow ij$ and cross section $\langle\sigma_{ab}|v|\rangle$ for scattering $ab\rightarrow ij$ as[25]
@@
where $n_{a}^{\rm eq}$ is the equilibrium number density of a particle $a$. The present DM abundance can be expressed as $\Omega_{DM}=M_{DM}Y_{DM}^{\infty}s_{0}/\rho_{0}$, where $\rho_{0}$ and $s_{0}$ are the energy density and the entropy density in the present Universe. They are given as $\rho_{0}=3M_{\rm pl}^{2}H_{0}^{2}$ and $s_{0}=2.27\times 10^{-38}$ GeV3, and then it determines the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ through $\Omega_{N_{1}}h^{2}=0.12$ as $Y_{N_{1}}^{\infty}=4.2\times 10^{-10}(M_{1}/{\rm GeV})^{-1}$.
Here, we roughly estimate both the freezeout temperature $T_{\eta}$ of the coannihilation of $\eta_{R}$ and a value of $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$. The freezeout temperature $T_{\eta}(\equiv M_{\eta_{R}}/z_{\eta})$ can be estimated through a condition $2n_{\eta_{R}}^{\rm eq}(T_{\eta})\langle\sigma_{\eta\eta}|v|\rangle=H(T_{\eta})z_{\eta}$ by assuming instantaneous freezeout as found from eq. (15). Using the Hubble parameter $H(T)$ given in eq. (19), $z_{\eta}$ can be determined by solving this condition as [26]
$z_{\eta}=\ln\frac{0.384M_{\rm pl}\langle\sigma_{\eta\eta}|v|\rangle_{\eta}M_{\eta_{R}}}{(g_{\ast}z_{\eta})^{1/2}}.$ (18)
The $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$ can be estimated from the condition $\Gamma_{\eta_{R}}=H(T_{\eta})$. The $\eta_{R}$ decay width $\Gamma_{\eta_{R}}$ and the Hubble parameter $H(T)$ are given as
@@
where $c_{p}$ is a constant determined by the phase space volume.cccWhen the mass of $\eta_{R}$ and $N_{1}$ take closely degenerate values, $c_{p}$ can be represented as $c_{p}=2\delta_{1}$, where $\delta_{1}$ is defined by $M_{1}=(1-\delta_{1})M_{\eta_{R}}$ with $\delta_{1}\ll 1$.
@@
If $h_{1}$ takes a larger value than the one given in eq. (20), the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be approximated well as $Y_{N_{1}}^{\infty}=Y_{\eta_{R}}^{\rm eq}(T_{c})$ through the freezein caused by the $\eta_{R}$ decay, where $T_{c}$ is defined by $Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$ since the $\eta_{R}$ decay at $T<T_{c}$ does not affect $Y_{N_{1}}^{\infty}$ substantially. $Y_{N_{1}}^{\infty}$ is determined by the coupling $h_{1}$ for a fixed $M_{1}$ and the freezeout of the $\eta_{R}$ (co)annihilation does not affect the result. However, this situation could change when $h_{1}$ takes a smaller value than the one given in eq. (20) and $T_{c}<T_{\eta}$ is satisfied. In this case, the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be affected by the equilibrium number density of $\eta_{R}$ at $T_{\eta}$. $T_{\eta}$ has been known to be determined through eq. (18) by the couplings $\lambda_{3}$ and $\lambda_{4}$ through the study of the inert doublet DM model [24]. If we use this knowledge, the required $N_{1}$ relic abundance could be predicted through a relation $Y_{N_{1}}^{\infty}\simeq Y_{\eta_{R}}^{\rm eq}(T_{\eta})$. It also gives constraints on the couplings $\lambda_{3}$ and $\lambda_{4}$.

### v2

@@ Abstract
A right-handed neutrino is a promising candidate for dark matter (DM) which has no interaction with nuclei. Since two right-handed neutrinos explain neutrino oscillation data and baryon number asymmetry through both the seesaw mechanism and leptogenesis based on the out-of-equilibrium decay of the lighter one, the lightest third right-handed neutrino could be free from them and stable to be DM. It suggests a possibility that the right-handed neutrinos could be common key particles for the explanation of three problems in the standard model (SM). We study this possibility by analyzing the structure of a neutrino mass matrix in the scotogenic model for neutrino mass.
@@ 1 Introduction
Neutrino oscillation experiments prove the existence of nonzero neutrino mass [1]. Observations of both the abundances of light elements and the cosmic microwave background suggest the existence of baryon number asymmetry in the Universe [2, 3]. However, both the neutrino mass and the baryon number asymmetry cannot be explained in the standard model (SM). It suggests that the SM is incomplete and has to be extended. If the SM is extended by two right-handed neutrinos which are assumed to have Majorana mass, these two problems are considered to be solved through the so called seesaw mechanism [4] and leptogenesis caused by the out-of-equilibrium decay of the right-handed neutrino [5] as long as their Majorana mass is larger than $O(10^{9})$ GeV [6].
On the other hand, the SM has no candidate of dark matter (DM) whose existence has been proved through various astrophysical and cosmological observations. Its direct detection experiments severely constrain its mass and elastic scattering cross section with a nucleus. As a result, a lot of DM candidates are excluded from their consistency with the required DM relic abundance [7]. It may suggest a possibility that DM has no interaction with nuclei. In this sense, the right-handed neutrino, which is massive and electrically neutral, could be a good candidate of DM since it has no interaction with nuclei. A possibility such that a right-handed neutrino is DM has been studied from a various view points. One may find recent works in [8, 9, 10, 11]. If we combine this idea with the above solution for the neutrino mass and the baryon number asymmetry, the third right-handed neutrino seems to be a good DM candidate if its stability is guaranteed by some symmetry. This requirement for the symmetry may impose to introduce an additional doublet scalar $\eta$ since we have to guarantee the existence of their invariant Yukawa couplings between doublet leptons and right-handed neutrinos under the assumed symmetry. As such an example, we have the scotogenic model [12], which can make the neutrino mass generation possible radiatively around TeV scales. The model seems to be very interesting as a simple extension of the SM since three big problems in the SM might be explained by the right-handed neutrinos simultaneously.
@@
In previous studies aiming to solve these problems including the case where DM is the lightest neutral component of $\eta$, the model is extended by some way, for example, new interaction is introduced to the right-handed neutrinos [15, 16, 17, 18]. In this paper, we propose a solution to these problems within the original scotogenic model without considering any extension of the model.aaaIt is discussed that $N_{1}$ can be DM consistently with the LFV constraint by considering coannihilation among the right-handed neutrinos without any extension [19].
The remaining parts are organized as follows. In section 2, after we review the model briefly, we discuss a neutrino mass matrix derived from it in detail, and flavor mixing and mass eigenvalues are fixed. Using them, we study a solution consistent with both the relic DM abundance and the constraint from the lepton flavor violating processes. In section 3, we discuss the consistency of the scenario with successful leptogenesis. We summarize the paper in section 4.
@@ 2 Neutrino model with right-handed neutrino DM
for the vacuum stability. If $\eta$ is assumed to have no vacuum expectation value (VEV), neutrino mass is not produced at tree level and the $Z_{2}$ symmetry is kept as an exact symmetry of the model.
@@
where $\eta^{\pm}$ stands for charged components. $\eta_{R}$ and $\eta_{I}$ are real and imaginary part of the neutral component, respectively. In this setting, a lightest right-handed neutrino $N_{1}$ or $\eta_{R}$ can be stable due to the $Z_{2}$ symmetry. We suppose that $N_{1}$ is DM and the baryon number asymmetry is generated by leptogenesis through the out-of-equilibrium decay of a heavier right-handed neutrino. In the following study, mass ordering of these extra particles is assumed to satisfy [20]
@@ 2.1 Neutrino mass and flavor mixing
Although neutrino mass is forbidden at tree level in Lagrangian (1), small neutrino mass is induced through a one-loop diagram caused by the Higgs VEV $\langle\phi\rangle$ as
@@
Its diagonalization matrix can be described by using the Yukawa couplings $h_{\alpha k}$ explicitly. When a condition $\sum_{\alpha}(h_{\alpha j}h_{\alpha k})\propto\delta_{jk}$ is satisfied, which corresponds to the condition that matrices $\mu^{j}$ and $\mu^{k}$ in eq. (5) are commutable [15], ${\cal M}_{\nu}$ is diagonalized by a matrix $U$ as $U^{T}{\cal M}_{\nu}U={\rm diag}(m_{1},m_{2},m_{3})$. This $U$ can be expressed as
@@
where $h_{k}$ is a real constant. An absolute value of the mass eigenvalue $m_{k}$ is fixed as
@@
where $p_{k}$ is defined as $p_{k}=\left|\sum_{\alpha}U_{\alpha k}^{2}\right|$. Here, it should be noted that $\sum_{\alpha}U_{\alpha k}^{2}=1$ is not satisfied if the Yukawa coupling $h_{\alpha k}$ is complex. By using this $U$, the PMNS matrix [21] can be written as $V_{\rm PMNS}=U_{e}^{\dagger}UP$ , where $U_{e}$ is a matrix which is used to diagonalize a charged lepton mass matrix and $P$ is a matrix of Majorana phases ${\rm diag}(1,e^{i(\gamma_{2}-\gamma_{1})},e^{i(\gamma_{3}-\gamma_{1})})$. Although $V_{\rm PMNS}$ is unitary in a case where $h_{\alpha k}$ is real, it is not unitary in a case where $h_{\alpha k}$ is complex since $U$ is not a unitary matrix.
@@
$U=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}&c_{12}c_{23}-s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}&-c_{12}s_{23}-s_{12}s_{13}c_{23}&c_{13}c_{23}\\ \end{array}\right),$ (8)
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}e^{i\alpha_{ij}}$. We can represent the neutrino Yukawa coupling $h_{\alpha k}$ and the mass eigenvalue $m_{k}$ by using this $U$ as eq. (6) and eq. (7), respectively.
If $h_{\alpha 1}$ has the same structure as $h_{\alpha 2}$ for example, the mass matrix ${\cal M}_{\nu}$ can be diagonalized by $U$ given in eq. (8) but it has a zero mass eigenvalue $m_{1}=0$. On the formula of $m_{2}$ in eq. (7), $p_{2}h_{2}^{2}\Lambda_{2}$ is replaced by $p_{1}h_{1}^{2}\Lambda_{1}+p_{2}h_{2}^{2}\Lambda_{2}$. As such a concrete example, neutrino Yukawa couplings may be taken as [18]
@@
In that case, $U$ is found to be the tribimaximal mixing matrix [22] and the mixing angles are fixed as
@@
The fault of the tribimaximal mixing matrix as the PMNS matrix could be overcome by considering a nontrivial flavor mixing matrix $U_{e}$ in the charged lepton sector [18]. Since $U_{e}$ is expected to be almost diagonal, it can become a phenomenologically viable possibility.
Here, we remark on an important consequence caused by the feature of this mass matrix addressed above. If we consider leptogenesis based on the out-of-equilibrium decay of the right-handed neutrino $N_{k}$, $CP$ asymmetry in this process is known to be caused by the interference between tree and one-loop diagrams and it should be nonzero at least to make leptogenesis work. On this point, the model predicts an interesting result. Since this $CP$ asymmetry is proportional to ${\rm Im}(\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j})^{2}$ for $k\not=j$ as shown in eqs. (21) and (25) later, this condition cannot be satisfied because of $\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j}\propto\delta_{kj}$ for the real neutrino Yukawa coupling $h_{\alpha k}$. This is the case even if Majorana phases in the right-handed neutrino mass are taken into account. Thus, the Yukawa coupling $h_{\alpha k}$ should be complex, and the PMNS matrix cannot be unitary in that case. However, we should note that there is an exceptional case where the $CP$ asymmetry can take a nonzero value even if the Yukawa coupling $h_{\alpha k}$ is real. The existence of a zero mass eigenvalue presents such a case as found in an example given in eq. (9) when Majorana phases exist in the right-handed neutrino mass [18]. The PMNS matrix becomes unitary in this case.
@@
If we use the best fit values from neutrino oscillation data given in [23] for $\Delta m_{ij}^{2}$ and $\sin\theta_{ij}$ in the matrix $U$, we can calculate two effective masses relevant to two observables, that is, the effective masses for the neutrinoless double $\beta$ decay and the $\beta$ decay. They are predicted to be
$m_{\beta\beta}\simeq\left|\sum_{j=1}^{3}U_{ej}^{2}m_{j}\right|=0.0037~{\rm eV},\quad m_{\beta}\simeq\sqrt{\sum_{j=1}^{3}|U_{ej}|^{2}m_{j}^{2}}=0.0088~{\rm eV}$ (13)
in normal ordering case. These results are much smaller than their present bounds [24, 25].
@@
$M_{\eta_{R}}\simeq M_{2}\simeq M_{3}\simeq 3~{\rm TeV},\quad\lambda_{5}=-10^{-3},$
$h_{2}=1.11\times 10^{-3},\quad h_{3}=4.15\times 10^{-3},$ (14)
where we assume $\alpha_{12}=\alpha_{13}=0,~\alpha_{23}=\frac{\pi}{2}$ and use for $\sin\theta_{ij}$ the best fit values from neutrino oscillation data. $M_{1}$ and $h_{1}$ can be fixed freely as long as $h_{1}$ is small enough. We note that the present bounds of LFV like $\mu\rightarrow e\gamma$ [23] can be satisfied for these parameters. We come back to this point in the last part of the paper again.
@@ 2.2 Relic abundance of right-handed neutrino DM
Right-handed neutrino $N_{1}$ is stable to be DM under the condition (4). However, its abundance is difficult to reach the required value in the freezeout scenario as mentioned in the introduction. In fact, it can be produced in the thermal bath at the temperature higher than its mass through both the $\eta$ decay and the scatterings ($\eta^{\dagger}\eta\rightarrow N_{1}N_{1}^{c}$ and $\ell_{\alpha}^{\dagger}\ell_{\alpha}\rightarrow N_{1}N_{1}^{c}$) caused by the neutrino Yukawa coupling $h_{\alpha 1}$ if it is large enough. In that case, $h_{\alpha 1}$ has to take a value of $O(1)$ to reduce its abundance sufficiently through the annihilation. Unfortunately, such a large Yukawa coupling causes LFV processes like $\mu\rightarrow e\gamma$ at an unacceptable level through the one-loop diagram [13]. Since this difficulty is caused by a feature that $N_{1}$ has interaction due to the neutrino Yukawa couplings only, new interaction like $U(1)^{\prime}$ has been introduced to reduce its abundance in many articles [15, 16].
In the present model, since the neutrino Yukawa coupling $h_{1}$ is supposed to be very small as shown in eq. (14), its thermal abundance is originally difficult to reach a value larger than the required one so that the freezeout scenario cannot be applied to the model. However, even in that case, we can consider the freezein scenario [26] as an alternative possibilitydddA right-handed neutrino DM through the freezein mechanism has been studied in the different model [8] and in the original scotogenic model [10, 11]. in which $N_{1}$ is produced through the $\eta$ decay and the scatterings caused by the neutrino Yukawa coupling $h_{1}$. Since $h_{1}$ is supposed to be sufficiently small here, the $N_{1}$ generation through these processes is expected to be largely suppressed. On the other hand, the next lightest DM candidate $\eta_{R}$ is in the thermal equilibrium and decrease of its number density is controlled by its coannihilation [27] caused by the SM interactions. After the freezeout of this coannihilation, the relic $\eta_{R}$ finally decays to $N_{1}$. Thus, we have to take into account that there is a case where the contribution from such $N_{1}$ plays a crucial role in the estimation of the $N_{1}$ relic abundance.
@@
where $H(M_{2})$ represents the Hubble parameter at $T=M_{2}$. Reaction densities $\gamma^{D}(a)$ and $\gamma^{S}(ab)$ are respectively related to thermally averaged decay width $\langle\Gamma_{a}\rangle$ for $a\rightarrow ij$ and thermally averaged cross section $\langle\sigma_{ab}|v|\rangle$ for the scattering $ab\rightarrow ij$ as[29]
@@
where $n_{a}^{\rm eq}$ is the equilibrium number density of a particle $a$. The present DM abundance can be expressed as $\Omega_{DM}=M_{DM}Y_{DM}^{\infty}s_{0}/\rho_{0}$, where $\rho_{0}$ and $s_{0}$ are the energy density and the entropy density in the present Universe. Since they are given as $\rho_{0}=3M_{\rm pl}^{2}H_{0}^{2}$ and $s_{0}=2.27\times 10^{-38}$ GeV3, the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is found to be expressed as $Y_{N_{1}}^{\infty}=4.2\times 10^{-10}(M_{1}/{\rm GeV})^{-1}$ from $\Omega_{N_{1}}h^{2}=0.12$.
Here, we roughly estimate both the freezeout temperature $T_{\eta}$ of the coannihilation of $\eta_{R}$ and a value of $h_{1}$ which makes the $\eta_{R}$ decay reach the equilibrium at $T_{\eta}$. Such a value of $h_{1}$ can be estimated from the condition $\Gamma_{\eta_{R}}=H(T_{\eta})$. The $\eta_{R}$ decay width $\Gamma_{\eta_{R}}$ and the Hubble parameter $H(T)$ are given as
@@
where $c_{p}$ is a constant determined by the phase space volume.eeeWhen the masses of $\eta_{R}$ and $N_{1}$ take closely degenerate values, $c_{p}$ can be represented as $c_{p}=2\delta_{1}$ where $\delta_{1}$ is defined by $M_{1}=(1-\delta_{1})M_{\eta_{R}}$ with $\delta_{1}\ll 1$. The freezeout temperature $T_{\eta}(\equiv M_{\eta_{R}}/z_{\eta})$ can be estimated through a condition $2n_{\eta_{R}}^{\rm eq}(T_{\eta})\langle\sigma_{\eta\eta}|v|\rangle=H(T_{\eta})z_{\eta}$ as found from eq. (15) by assuming instantaneous freezeout. Using the Hubble parameter $H(T)$ given in eq. (18), $z_{\eta}$ can be determined by solving this condition as [30]
$z_{\eta}=\ln\frac{0.384M_{\rm pl}\langle\sigma_{\eta\eta}|v|\rangle M_{\eta_{R}}}{(g_{\ast}z_{\eta})^{1/2}}.$ (20)
@@
If $h_{1}$ takes a larger value than the one given in eq. (19), the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be approximated well as $Y_{N_{1}}^{\infty}=Y_{\eta_{R}}^{\rm eq}(T_{c})$ through the freezein caused by the $\eta_{R}$ decay. $T_{c}$ is defined by $Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$ since the $\eta_{R}$ decay at $T<T_{c}$ does not affect $Y_{N_{1}}^{\infty}$ substantially. $Y_{N_{1}}^{\infty}$ is determined only by the coupling $h_{1}$ for a fixed $M_{1}$ and the freezeout of the $\eta_{R}$ coannihilation does not affect this result. However, this situation could change when $h_{1}$ takes a smaller value than the one given in eq. (19) and $T_{c}<T_{\eta}$ is satisfied. In this case, the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be affected by the equilibrium number density of $\eta_{R}$ at $T_{\eta}$ since $\eta_{R}$ decays to $N_{1}$ finally. $T_{\eta}$ is determined through eq. (20) by $\langle\sigma_{\eta\eta}|v|\rangle$ which depends on the couplings $\lambda_{3}$ and $\lambda_{4}$. In the study of the inert doublet DM model [28], these couplings are known to play a crucial role there. If we take account of it, the $N_{1}$ relic abundance could be predicted through a relation $Y_{N_{1}}^{\infty}\simeq Y_{\eta_{R}}^{\rm eq}(T_{\eta})$. It also gives constraints on the couplings $\lambda_{3}$ and $\lambda_{4}$.